Use the function rule to complete the table. -10x+y=4
x -2, -1, 0, 1, 2
y __, __, __, __, __

Please show work.

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Use the function rule to complete the table. -10x+y=4
x -2, -1, 0, 1, 2
y __, __, __, __, __
=============================
y = 10x + 4
f(-2) = 10(-2)+4 = -16
f(-1) = 10(-1)+4 = -6
f(0) = 4
f(1) = 14
f(2) = 24
------------
Cheers,
Stan H.
----------------

Answer by solver91311(24713)



Solve the equation for in terms of everything else by adding to both sides, thus:



On your paper, write the equation in pencil. For the value -2, erase the and write in -2 in its place. Do the arithmetic to get the value of that corresponds to the -value -2. Repeat the process for each of the other given values for

John

My calculator said it, I believe it, that settles it